an equation in point- slope form and slope- intercept form for (Passing through (-2,-7) and parallel to the line whose equation is y=-5x+4)
Answer by lwsshak3(11628)

an equation in point- slope form and slope- intercept form for (Passing through (-2,-7) and parallel to the line whose equation is y=-5x+4
***
given equation: y=-5x+4
slope=-5
..
slope-intercept form:y=mx+b, m=slope, b=y-intercept
equation of parallel line:
y=-5x+b
solve for b, using coordinates of given point(-2,-7)
-7=-5*(-2)+b
b=-17
equation: y=-5x-17
5x+y=-17
..
point-slope form:(y-y1)=m(x-x1)
y-(-7)=-5(x-(-2))
y+7=-5(x+2)
y+7=-5x-10
5x+y=-17